OpenStudy (anonymous):

I found the solutions ( think they are right) but I am having trouble finding the x intercepts. The solutions I got are -sqrt(2), sqrt(2). The problem is 3x^2-6.

OpenStudy (amistre64):

x intercepts occur when y=0

OpenStudy (amistre64):

3x^2 -6 = 0 3x^2 = 6 x^2 = 2 so yeah, +- sqrt(2)

OpenStudy (anonymous):

so my solutions are also my x intercepts?

OpenStudy (amistre64):

you really didnt provide something that was "solvable". it was simply an expression how did you get your "solutions"?

OpenStudy (anonymous):

I actually called a friend who tried to help me work thru it but they didn't understand it either. Thats what we come up with. So I dunno.

OpenStudy (amistre64):

the set of ordered pairs that would fit the expression draw a parabola on a graph

OpenStudy (amistre64):

the places where that graph hits the y=0 (x ints) are then +- sqrt(2) when we solve for f(x) = 0

OpenStudy (anonymous):

Ok i will try that...thank you.